You need a total of 500,000 grams of fuel to make it back home. If you can produce 6.02X1025 molecules of H2 per hour, how many days will it take to refuel?

In the time it takes for your ship to refuel, how many liters of oxygen gas will you inhale to survive?
1
Expert's answer
2019-01-22T04:35:55-0500

1. Find how many molecules of H2 are in the 500,000 grams of fuel:

n=m/M

M(H2) = 1.01*2 = 2.02 g/mol

n(H2) = 500,000 g/2.02 = 247,525 mol

N(H2) = n*NA = 247,525 *6.022*1023 = 1.49*1029

Find how many hours will it take to produce 1.49*1029 molecules of H2

T(hours) = 1.49*1029/6.02*1025 = 2475.25 hours

As thre are 24 hours in a day then

T(days) = 2475.25/24 = 103.14 days

2. An average human needs 785.7 g of pure oxygen per day.

Find mass of oxygen consumed by a person during 103.14 days:

m(O2) = 103.14 days*785.7 g/day = 81037.1 g

n= m/M

M(O2) = 16.00*2 = 32.00 g/mol

n(O2) = 81037.1 g /32.00 g/mol = 2532.4 mol

At STP molar volume Vm = 22.4 L/mol

V= Vm *n = 22.4 L/mol *2532.4 mol = 56726 L

Answer:

1. 103.14 days

2. 56726 L